OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. 4 - 24 + 144 - 864 + ... summation of four times six to the power of n from n equals zero to infinity summation of four times six to the power of n from n equals zero to infinity summation of four times negative six to the power of the quantity n minus one from n equals zero to infinity summation of four times six to the power of the quantity n plus one from n equals zero to infinity

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

\[\sum_{n=D}^{\infty} 4 X 6n\]

OpenStudy (anonymous):

4 times^ six n

OpenStudy (aravindg):

Make sure you check the sign too!

OpenStudy (anonymous):

Here are the answer responses:

OpenStudy (anonymous):

would it be the last one? \[\sum_{n=D}^{\infty} 4 \times 6^{n}\]

OpenStudy (aravindg):

The last option always give a positive value for all n. But we also get negative values alternatively in the sequence. Hence it is wrong.

OpenStudy (anonymous):

Ohhh okay, can you explain how to solve it? I'm so confused on this

OpenStudy (aravindg):

Actually you are 99% complete. You just do a manipulation on ur answer so that u get alternative +-

OpenStudy (anonymous):

Ohh! so its this one?

OpenStudy (aravindg):

Exactly! :)

OpenStudy (anonymous):

Thankyou so much! :) you were an awesome helper!

OpenStudy (aravindg):

yw :)